OpenStudy (anonymous):

can someone help me with this? lim(2x+1) / (x-2), x ->infinity thank you

OpenStudy (anonymous):

\[\lim_{x\to\infty}\frac{2x+1}{x-2}\] Divide through by the highest power of \(x\): \[\displaystyle\lim_{x\to\infty}\dfrac{2+\dfrac{1}{x}}{1-\dfrac{2}{x}}\] What happens to the fractions as \(x\to\infty\)?

OpenStudy (anonymous):

tends to 2. Thank you! I have too much to learn. 8)

OpenStudy (anonymous):

tends to 0, I mean. Thank you.